[Brahms-l] Shift report 20040112 00:00-08:00

From: Apache <apache@bnl.gov>
Date: Mon Jan 12 2004 - 07:56:01 EST
------------------------------------------------------------
Date: Mon Jan 12 07:56:00 2004, Shift: 00:00-08:00
Supervisor: CEJ, with: 
Log book#15 pages 30 - 30

Shift Summary:
------------------------------------------------------------
1) Nice store from 1.30. Gap cleaning started 6.15. 
   Beam dumped 6.45.   

2) pedestal runs: 9559, 9560, 9570, 9571
   FS @ 4deg D1 = B3450, MRS @ 90 deg D5=A1050
   run     nKtr3  nKtr6 
   9562 :    97    300
   9563 :    66    200 
   9564 :    74    200
   9565 :    50    200
   9566 :    27    100  
   total: ~ 300   1000
   FS @ 4deg D1 = B2442, MRS @ 90 deg D5=A1050
   9567 :    50    200
   9568 :    48    200
   9569 :     8     50
   total: ~ 100    450

3) T5-1 X1,X2,U1, U2 planes have dead wires.
   The beam condition in E-log (when new run is 
   started) is from jan 7.
   The D4 WFG loopback seems to be stuck - the 
   measured value is good.
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Mon Jan 12 07:56:20 2004

This archive was generated by hypermail 2.1.8 : Mon Jan 12 2004 - 07:56:38 EST